Dynamite Industries paid a dividend of RM1.65 for its common stock yesterday. The
dividends of company are expected to grow at 9% per year indefinitely. If the risk free
rate is 3% and investors' risk premium on this stock is 8%, what is the estimate value
of Dynamite Industries stock 2 years from now?
1
Expert's answer
2015-03-20T10:54:39-0400
Dynamite Industries paid a dividend of div0 = RM1.65, g = 9% per year indefinitely, risk free
rate is 3% and investors' risk premium on this stock is 8%.
The formula for the present value of a stock with constant growth is the
estimated dividends to be paid divided by the difference between the
required rate of return and the growth rate.
Required rate of return is:
r = 0.03 + 2*(0.08 - 0.03) = 0.13
The estimate value of Dynamite Industries stock 2 years from now will be:
P = div0/(r - g) = 1.65/(0.13 - 0.09) = RM41.25.